Write an equation in slope-­intercept form for the line that satisfies each set of conditions:
1. passes through (­-6, 15), parallel to the graph of 3x + 2y = 1
2. passes through (5, ­-2), perpendicular to the graph of x + 2y = 8

Answer by Alan3354(69443)

Write an equation in slope-­intercept form for the line that satisfies each set of conditions:
1. passes through (­-6, 15), parallel to the graph of 3x + 2y = 1
Find the slope of the given line.
To do that, put it in slope-intercept form. (That means solve for y)
3x + 2y = 1
2y = -3x + 1
y = (-3/2)x + 1/2
The slope, m is the coefficient of x, -3/2
----
Parallel lines have the same slope.
Use y-y1 = m*(x-x1) where (x1,y1) is the point (-6,15).
y-15 = (-3/2)*(x+6)
y-15 = (-3/2)x - 9
y = (-3/2)x + 6
=======================
2. passes through (5, ­-2), perpendicular to the graph of x + 2y = 8
Find the slope, as above.
y = (-1/2)X + 4
The slope of perpendicular lines is the negative inverse, --> m = 2.
From here, same as #1.